How many milliliters of water have to be added to 300 mL of stock solution of 11.1 M HNO3 to have final concentration of 6.00 M?
1
Expert's answer
2019-11-26T06:39:50-0500



Task:

How many milliliters of water have to be added to 300 mL of stock solution of 11.1 M HNO3 to have final concentration of 6.00 M?


Solution:

This dilution problem uses the equation

С1V1 = C2V2;


C1 = 11.1 M - the initial molarity (concentration)

V1 = 300 mL - the initial volume

C2 = 6 M - the desired molarity (concentration)

V2 = (300 + x mL) - the volume of the desired solution


(11.1 M) (300 mL) = (6.00 M)(300 mL + x )

3330 M mL= 1800 M mL + 6.00x M

(3330-1800) = 6.00x

1530 = 6.00x

1530 / 6.00 = x

x = 255

255 mL needs to be added to the original 300 mL solution in order to dilute it from 11.1 M to 6.00 M.


Answer: 255 mL of water have to be added.